OpenStudy (anonymous):

If 7,500 bacteria, with a growth constant k=0.4, are present 3hrs after beginning the experiment, how many are present 8 hrs later? a) 24,849 bacteria b) 2,259 bacteria c) 610,500 bacteria d) 184,000 bacteria

OpenStudy (amistre64):

i assume the k refers to 'percentage'

OpenStudy (amistre64):

7500 = P(1.4)^3

OpenStudy (amistre64):

P= 7500/(1.4)^3 X = [7500/(1.4)^3](1.4)^11

OpenStudy (amistre64):

the closest number i get to the choices is d, but i cant say for sure that I was right

OpenStudy (amistre64):

i see it; e^kt ....

OpenStudy (amistre64):

7500 = P.e^(.4*3)

OpenStudy (amistre64):

183 ....... so its still D but with more confidence lol

OpenStudy (anonymous):

okay thank you